OpenStudy (anonymous):

The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used. B = 137°, c = 6, b = 11

OpenStudy (danjs):

sin137/ 11 = sin(C)/6 then angle A = 180-angle B - angle C

OpenStudy (anonymous):

213.73 ft^3?

OpenStudy (danjs):

and same again for side A

OpenStudy (anonymous):

im stuck

OpenStudy (anonymous):

theres 2 choices with a as 5.82

OpenStudy (trojanpoem):

Write the choices.

OpenStudy (anonymous):

The triangle cannot be solved with the Law of Sines. C = 21.2°, A = 21.8°, a ≈ 5.8 No triangle is formed. C = 21.8°, A = 21.2°, a ≈ 5.8

OpenStudy (anonymous):

its either the 2nd or 4th

OpenStudy (trojanpoem):

angle a = 21.16 ~~ 21.2

OpenStudy (trojanpoem):

The sine law state that: asina=bsinb=csinc=2r=d 11sin137=6sinc <c = 21.84 From (1) ; <a = 21.16 asin21.16=11sin137 a = 5.82 a > c ( no triangle can be formed )

OpenStudy (trojanpoem):

This one : No triangle is formed. C = 21.8°, A = 21.2°, a ≈ 5.8

OpenStudy (anonymous):

thanks man!

OpenStudy (trojanpoem):

Any time.